# Synergistic Dynamical Decoupling and Circuit Design for Enhanced Algorithm Performance on Near-Term Quantum Devices

### v1

@@ Abstract
Dynamical decoupling (DD) is a promising technique for mitigating errors in near term quantum devices. However, its effectiveness depends on both hardware characteristics and algorithm implementation details. This paper explores the synergistic effects of dynamical decoupling and optimized circuit design in maximizing the performance and robustness of algorithms on near term quantum devices. By utilizing eight IBM quantum devices, we analyze how hardware features and algorithm design impact the effectiveness of DD for error mitigation. Our analysis takes into account factors such as circuit fidelity, scheduling duration, and hardware native gate set. We also examine the influence of algorithmic implementation details including specific gate decompositions, DD sequences, and optimization levels. The results reveal an inverse relationship between the effectiveness of DD and the inherent performance of the algorithm. Furthermore, we emphasize the importance of gate directionality and circuit symmetry in improving performance. This study offers valuable insights for optimizing DD protocols and circuit designs, highlighting the significance of a holistic approach that leverages both hardware features and algorithm design for high quality and reliable execution of near term quantum algorithms.
@@ I Introduction
Near-term quantum devices (NISQ) [1] hold immense potential but face hurdles in accuracy and reliability due to inherent noise arising from environmental fluctuations, imperfect gate operations, and qubit interactions. Moreover, limitations in qubit count and connectivity restrict the complexity of achievable quantum circuits. Robust error mitigation techniques [2] are therefore crucial to unlock the full potential of NISQ devices. Dynamical decoupling (DD) [3, 4, 5, 6] stands out as a powerful approach for NISQ devices due to its simplicity and low resource overhead. It mitigates decoherence errors by applying a carefully designed sequence of control pulses during idle periods of the qubits. These pulses effectively suppress the unwanted interaction between qubits and their environment, protecting the desired quantum state. DD has been demonstrated in various quantum systems including spins [7, 8, 9, 10, 11, 12], superconducting qubits [5, 13, 14], and trapped ions [15]. The effectiveness of DD extends beyond decoherence suppression, as it can also mitigate crosstalk [13, 16] and coherent errors [17].
@@
Numerous DD sequences have been developed, with prominent examples including the Carr-Purcell (CP) [18], Carr-Purcell-Meiboom-Gill (CPMG) [19], XY4 [20, 21, 22, 23], KDD [24], and Uhrig dynamical decoupling (UDD) [25]. However, the effectiveness of different sequences varies significantly [26]. Prior work has shown that the CPMG sequence outperforms the CP sequence [4]. Additionally, higher-order sequences generally outperform lower order sequences [4, 26], but optimizing pulse intervals within CPMG and XY4 sequences can achieve comparable performance [26]. In the context of the quantum approximate optimization algorithm (QAOA) [27, 28, 29, 30, 31], DD sequences like CP, CPMG, and XY4 significantly enhance performance, while KDD showed worse performance on IBM quantum processing units (QPUs) [32]. Combining DD sequences with pulse-level optimization in QAOA can further improve performance [32]. While DD sequences mitigate errors, these additional control pulses can also introduce errors such as gate infidelities or crosstalk. To address this trade-off, adaptive approaches have been explored [33], which estimate the potential benefit of DD for each qubit combination and selectively apply DD to the subset that offers the most benefit. Furthermore, recent work has explored the use of empirical learning schemes for enhancing DD effectiveness on quantum processors [34], leveraging data-driven techniques to optimize the implementation of DD sequences, leading to improved error suppression capabilities.
Despite extensive research on DD sequences, most studies neglect the crucial role of algorithm implementation details such as transpilation efficiency [35, 36, 37, 38], circuit structure, and gate decompositions [39]. This work addresses this gap by examining the factors influencing DD sequence effectiveness, emphasizing the critical role of algorithm implementation details in enhancing overall performance. We focus on the CPMG sequence due to its well-established effectiveness, robustness [4], and good performance for QAOA on IBM QPUs [32].
@@
The study explores how hardware characteristics and algorithm design affect the effectiveness of DD for error mitigation on near term quantum devices by analyzing different implementations of the QAOA on eight IBM QPUs. Our findings reveal an inverse relationship between the initial performance of the algorithm without DD and the effectiveness of DD. While DD generally enhances performance and robustness, circuits with inherently higher fidelity and shorter execution times benefit less from DD than those with lower initial performance. Moreover, factors such as hardware native gates of QPUs, chosen gate decomposition strategy, and optimization level can also influence the effectiveness of DD. Additionally, using gates with consistent directionality and maintaining circuit symmetry during design lead to improved performance. These findings emphasize the importance of a holistic approach that considers both hardware and software optimization for the successful execution of algorithms with DD on NISQ devices, providing valuable insights for optimizing DD protocols and designing more robust quantum algorithms.
@@
The paper is structured as follows. Section II outlines the methodology, presenting the benchmark circuits and metrics used to evaluate algorithm performance. Hardware considerations and a proposed synergistic design approach combining both hardware and algorithmic factors are also described. Section III then analyzes results, exploring the impact of various hardware factors including circuit fidelity, schedule duration, and native gate sets, as well as algorithmic factors including different implementations, DD sequences, and optimization levels. Finally, Sec. IV discusses the key findings and concludes.
@@ II Methodology
We experimentally investigate the impact of hardware and algorithmic factors on the effectiveness of dynamical decoupling in superconducting quantum processes. Hardware factors, such as circuit fidelity, schedule duration, and native gate sets, define the fundamental capabilities of a quantum device. Conversely, algorithmic factors encompass the design choices made during algorithm implementation (specific sequence of quantum operations), error suppression strategy (selection of a DD sequence), and circuit optimization techniques. These algorithmic choices ultimately determine how efficiently the inherent capabilities of the hardware are utilized for optimal performance. Understanding these factors can potentially improve the practical implementation of algorithms on near term quantum devices.
@@ II.1.1 QAOA for portfolio optimization
We employ QAOA for portfolio optimization as benchmarks. The portfolio optimization problem aims to select the optimal portfolio from all available options to maximize expected returns while minimizing financial risk. The QAOA has the potential to address this challenge [40, 41, 42]. The cost Hamiltonian, which describes the portfolio optimization problem for $n$ available assets, is expressed as [41]
$\hat{H}_{c}=\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}c_{i,j}\hat{Z}_{i}\hat{Z}_{j}-\sum_{i=1}^{n}k_{i}\hat{Z}_{i},$ (1)
where $c_{i,j}$ and $k_{i}$ are parameters specific to portfolio optimization, $\hat{Z}_{i}\hat{Z}_{j}$ and $\hat{Z}_{i}$ represent $ZZ$ interaction on qubits ($i$, $j$) and Pauli $Z$ operator acting on qubit $i$, respectively. The mixer Hamiltonian is given by [27]
@@
where $\hat{X}_{i}$ is Pauli $X$ operator acting on qubit $i$. After a QAOA depth of $p$, the total system evolves to |$\ket{\psi}=\prod_{j=1}^{p}e^{-i\gamma_{j}\hat{H}_{c}}e^{-i\beta_{j}\hat{H}_{m}}\ket{\psi_{0}}$, where |$\ket{\psi_{0}}$ is the eigenstate of the mixer Hamiltonian. The aim of QAOA is to find $2p$ parameters ($\beta_{1},...,\beta_{p},\gamma_{1},...,\gamma_{p}$) that minimize the expectation value of the cost Hamiltonian $F=\bra{\psi}\hat{H}_{c}\ket{\psi}$. We define the approximation ratio of QAOA as
@@
where $F_{0}$ represents the optimal value, and $F_{\max}$ signifies the worst-case value.
Our study uses QAOA with qubit numbers ranging from 3 to 12 and a depth of 1. To establish a baseline for experimental results, we present the simulation results conducted in a noise-free environment with Qiskit’s Qasm simulator in Table 1. All data use 30000 circuit repetitions (shots). We observe that approximation ratio and success probability decrease as the number of qubits increases.
Table 1: Noise-free simulation results of QAOA for portfolio optimization using Qiskit’s Qasm simulator with 30000 shots.
@@ II.1.2 Metric definition
where $r_{\epsilon}$ and $p_{\epsilon}$ represent approximation ratio and success probability obtained under noise conditions. $r_{0}$ and $p_{0}$ denote corresponding values obtained from the simulated noise free case (Table 1). For the same problem instances, we use identical values of $r_{0}$ and $p_{0}$ for evaluation. Due to noise, NAR and NSP typically exhibit values less than one. However, in rare instances, these metrics may exceed unity, indicating that specific noise patterns or randomness enhance performance compared to the simulated noise free case.
@@
Without any error mitigation being applied, NAR and NSP for algorithms executing on real noisy quantum devices are given by
$\mathrm{NAR}_{\mathrm{B}}$ $=$ $r_{b}/r_{0},$ (6)
$\mathrm{NSP}_{\mathrm{B}}$ $=$ $p_{b}/p_{0},$ (7)
@@
where B represents results obtained from real hardware without error mitigation. $r_{b}$ and $p_{b}$ represent the corresponding approximation ratio and success probability, respectively. When a specific error mitigation strategy, such as the DD sequence, is applied, NAR and NSP are given by
$\mathrm{NAR}_{\mathrm{DD}}$ $=$ $r_{d}/r_{0},$ (8)
$\mathrm{NSP}_{\mathrm{DD}}$ $=$ $p_{d}/p_{0},$ (9)
@@
where DD denotes the application of DD sequences. $r_{d}$ and $p_{d}$ represent approximation ratio and success probability, respectively, with error mitigation. To assess DD effectiveness in error mitigation, we introduce two additional metrics: $\Delta_{\mathrm{NAR}}$ and $\Delta_{\mathrm{NSP}}$. These metrics are defined as the difference between the corresponding values obtained with and without DD sequences:
$\Delta_{\mathrm{NAR}}$ $=$ $\mathrm{NAR}_{\mathrm{DD}}-\mathrm{NAR}_{\mathrm{B}},$ (10)
$\Delta_{\mathrm{NSP}}$ $=$ $\mathrm{NSP}_{\mathrm{DD}}-\mathrm{NSP}_{\mathrm{B}}.$ (11)
@@
A positive value for $\Delta_{\mathrm{NAR}}$ and $\Delta_{\mathrm{NSP}}$ indicates a successful improvement in performance resulting from the utilization of DD sequences.
@@
We further introduce the concept of error mitigation success rate (EMSR) to quantitatively assess the robustness of an error mitigation strategy. EMSR is defined as the percentage of experimental trials in which the application of error mitigation resulted in a better outcome compared to the unmitigated scenario. A high EMSR indicates a high success rate associated with the error mitigation protocol. A value of 100% suggests that applying this error mitigation strategy consistently improves algorithm performance across all experiments. In contrast, a low EMSR value suggests that the effectiveness of error mitigation may not be universally applicable and could potentially diminish performance. In this study, we quantify the robustness of DD error mitigation with two metrics: $\mathrm{EMSR}_{\mathrm{AR}}$ and $\mathrm{EMSR}_{\mathrm{SP}}$, depending on whether the approximation ratio or success probability is used. Any positive value of $\Delta_{\mathrm{NAR}}$ contributes to an increase in $\mathrm{EMSR}_{\mathrm{AR}}$, while any positive value of $\Delta_{\mathrm{NSP}}$ leads to an increase in $\mathrm{EMSR}_{\mathrm{SP}}$.
@@ II.2 Hardware considerations
This section provides information about IBM quantum devices used in our experiments. The QPUs with 27 qubits, namely ibmq_mumbai, ibmq_kolkata, ibm_cairo, and ibmq_ehningen, operate using basis gates {CX, ID, RZ, SX, X}, where ID represents identity gate, RZ performs a single qubit rotation around the $z$-axis, X is the NOT gate, and SX is the square root of X. On the other hand, the QPUs with 127 qubits, specifically ibm_kyoto, ibm_cusco, ibm_brisbane, and ibm_sherbrooke, utilize basis gates {ECR, ID, RZ, SX, X}, where ECR is the echoed cross resonance gate.
@@
Figure 1: Schedule of hardware native two-qubit gates: (a) CX gate on qubit pair (0,1) of ibm_cairo, and (b) ECR gate on qubit pair (0,1) of ibm_cusco. The system cycle time ($1\mathrm{dt}$) is 2/9 ns $\approx$ 0.22 ns in ibm_cairo, while it is 0.50 ns in ibm_cusco. $D_{i}$ represents the drive channel acting on qubit $i$, and $U_{j}$ is the control channel for a corresponding qubit pair $(c,t)$, driving the control qubit $c$ at the frequency of the target qubit $t$.
Figure 1(a) illustrates the schedule of a native CX gate on ibm_cairo. This implementation employs a single pulse gate duration of 112 dt and a cross resonance (CR) gate duration of 544 dt, leading to a total duration of 1312 dt, where dt represents the system cycle time. In contrast, Fig. 1(b) depicts the schedule of a native ECR gate on ibm_cusco. Here, the single-pulse and CR gate durations are 88 dt and 416 dt, respectively, resulting in a shorter total duration of 920 dt. The CX gate is typically implemented using one ECR gate and multiple single qubit gates. A further decomposition of CX gate into ECR and single pulse gates at pulse level enables the elimination of single pulse gates during circuit optimization [43]. Additionally, CX-based IBM QPUs, where CX is the native gate, support the operation of CX in two directions. On the other hand, ECR-based QPUs, where ECR is the native gate, typically only support the ECR gate in one direction. Consequently, quantum circuits designed for these latter devices need to be decomposed into sequences of gates that include only the supported direction of the ECR.

### v2

@@ Abstract
Dynamical decoupling (DD) is a promising technique for mitigating errors in near-term quantum devices. However, its effectiveness depends on both hardware characteristics and algorithm implementation details. This paper explores the synergistic effects of dynamical decoupling and optimized circuit design in maximizing the performance and robustness of algorithms on near-term quantum devices. By utilizing eight IBM quantum devices, we analyze how hardware features and algorithm design impact the effectiveness of DD for error mitigation. Our analysis takes into account factors such as circuit fidelity, scheduling duration, and hardware-native gate set. We also examine the influence of algorithmic implementation details, including specific gate decompositions, DD sequences, and optimization levels. The results reveal an inverse relationship between the effectiveness of DD and the inherent performance of the algorithm. Furthermore, we emphasize the importance of gate directionality and circuit symmetry in improving performance. This study offers valuable insights for optimizing DD protocols and circuit designs, highlighting the significance of a holistic approach that leverages both hardware features and algorithm design for the high-quality and reliable execution of near-term quantum algorithms.
@@ I Introduction
Near-term quantum (NISQ) devices [1] hold immense potential but face hurdles in accuracy and reliability due to inherent noise arising from environmental fluctuations, imperfect gate operations, and qubit interactions. Moreover, limitations in qubit count and connectivity restrict the complexity of achievable quantum circuits. Robust error mitigation techniques [2] are therefore crucial for unlocking the full potential of NISQ devices. Dynamical decoupling (DD) [3, 4, 5, 6] stands out as a powerful approach for NISQ devices due to its simplicity and low resource overhead. It mitigates decoherence errors by applying a carefully designed sequence of control pulses during idle periods of the qubits. These pulses effectively suppress the unwanted interaction between qubits and their environment, protecting the desired quantum state. DD has been demonstrated in various quantum systems, including spins [7, 8, 9, 10, 11, 12], superconducting qubits [5, 13, 14], and trapped ions [15]. The effectiveness of DD extends beyond decoherence suppression as it can also mitigate crosstalk [13, 16, 17, 18, 19] and coherent errors [20].
@@
Numerous DD sequences have been developed, with prominent examples including Carr–Purcell (CP) [21], Carr–Purcell–Meiboom–Gill (CPMG) [22], XY4 [23, 24, 25, 26], KDD [27], and Uhrig dynamical decoupling (UDD) [28]. However, the effectiveness of different sequences varies significantly [29]. Prior work has shown that the CPMG sequence outperforms the CP sequence [4]. Additionally, higher-order sequences generally outperform lower-order sequences [4, 29], but optimizing pulse intervals within CPMG and XY4 sequences can achieve comparable performance [29]. In the context of the quantum approximate optimization algorithm (QAOA) [30, 31, 32, 33, 34], DD sequences like CP, CPMG, and XY4 significantly enhance performance, while KDD shows worse performance on IBM quantum processing units (QPUs) [35]. Combining DD sequences with pulse-level optimization in QAOAs can further improve performance [35]. While DD sequences mitigate errors, these additional control pulses can also introduce errors such as gate infidelities or crosstalk. To address this trade-off, adaptive approaches have been explored [36], which estimate the potential benefit of DD for each qubit combination and selectively apply DD to the subset that offers the most benefit. Furthermore, recent work has explored the use of empirical learning schemes for enhancing DD effectiveness on quantum processors [37], leveraging data-driven techniques to optimize the implementation of DD sequences, leading to improved error suppression capabilities.
Despite significant research on DD sequences, the interplay between their effectiveness and specific algorithm implementations on real hardware remains underexplored. This study bridges this critical gap by investigating how factors such as transpilation efficiency [38, 39, 40, 41, 42], circuit structure [43, 44], and gate decompositions [45, 44] influence DD performance on IBM QPUs. By analyzing how these implementation details interact with the chosen DD sequence, we uncover synergistic effects that enhance overall algorithm performance on hardware. This focus on codesigning hardware and software offers deeper insights than separate studies of DD or individual circuit optimization techniques. We focus on the CPMG sequence due to its established effectiveness, robustness [4], and good performance for QAOAs on IBM QPUs [35].
@@
Our analysis of results from eight IBM QPUs demonstrates an inverse relationship between the initial performance of the algorithm without DD and the effectiveness of DD. While DD generally enhances performance and robustness, circuits with inherently higher fidelity and shorter execution times benefit less from DD than those with lower initial performance. Moreover, factors such as hardware-native gates of QPUs, the chosen gate decomposition strategy, and the optimization level can also influence the effectiveness of DD. Additionally, using gates with consistent directionality and maintaining circuit symmetry during design lead to improved performance. These findings emphasize the importance of a holistic approach that considers both hardware and software optimization for the successful execution of algorithms with DD on NISQ devices, providing valuable insights for optimizing DD protocols and designing more robust quantum algorithms.
@@
This paper is structured as follows. Section II outlines the methodology, presenting the benchmark circuits and metrics used to evaluate algorithm performance. Hardware considerations and a proposed synergistic design approach combining both hardware and algorithmic factors are also described. Section III then analyzes the results, exploring the impact of various hardware factors, including circuit fidelity, schedule duration, and native gate sets, as well as algorithmic factors, including different implementations, DD sequences, and optimization levels. Finally, Section IV discusses the key findings and concludes.
@@ II Methodology
We experimentally investigate the impact of hardware and algorithmic factors on the effectiveness of dynamical decoupling in superconducting quantum processes. Hardware factors, such as circuit fidelity, schedule duration, and native gate sets, define the fundamental capabilities of a quantum device. Conversely, algorithmic factors encompass the design choices made during algorithm implementation (specific sequence of quantum operations), error suppression strategy (selection of a DD sequence), and circuit optimization techniques. These algorithmic choices ultimately determine how efficiently the inherent capabilities of the hardware are utilized for optimal performance. Understanding these factors can potentially improve the practical implementation of algorithms on near-term quantum devices.
@@ II.1.1 QAOA for Portfolio Optimization
We employ QAOAs for portfolio optimization as benchmarks. The portfolio optimization problem aims to select the optimal portfolio from all available options to maximize expected returns while minimizing financial risk. The QAOA has the potential to address this challenge [46, 47, 48]. The cost Hamiltonian, which describes the portfolio optimization problem for $n$ available assets, is expressed as [47]
$\hat{H}_{c}=\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}c_{i,j}\hat{Z}_{i}\hat{Z}_{j}-\sum_{i=1}^{n}k_{i}\hat{Z}_{i}.$ (1)
The parameters $c_{i,j}$ and $k_{i}$ in Equation (1) depend on factors specific to the portfolio optimization problem, such as the covariance matrix and the return vector. In particular, $c_{i,j}=\frac{\lambda}{2}(q\sigma_{ij}+A)$ and $k_{i}=\frac{\lambda}{2}\left[A(2B-n)+(1-q)\mu_{i}-q\sum_{j=1}^{n}\sigma_{ij}\right]$, where $\lambda$ is the global scaling factor, $q$ is risk preference, $\sigma_{ij}$ is the covariance between assets $i$ and $j$, $A$ is the penalty factor, $B$ is the number of assets to be chosen, and $\mu_{i}$ is the expected return of asset $i$. The terms $\hat{Z}_{i}\hat{Z}_{j}$ and $\hat{Z}_{i}$ correspond to the $ZZ$ interaction on qubits ($i$, $j$) and Pauli $Z$ operator acting on qubit $i$, respectively. The mixer Hamiltonian is given by [30]
@@
where $\hat{X}_{i}$ is the Pauli $X$ operator acting on qubit $i$. After a QAOA depth of $p$, the total system evolves to |$\ket{\psi}=\prod_{j=1}^{p}e^{-i\gamma_{j}\hat{H}_{c}}e^{-i\beta_{j}\hat{H}_{m}}\ket{\psi_{0}}$, where |$\ket{\psi_{0}}$ is the eigenstate of the mixer Hamiltonian. The aim of the QAOA is to find $2p$ parameters ($\beta_{1},...,\beta_{p},\gamma_{1},...,\gamma_{p}$) that minimize the expectation value of the cost Hamiltonian $F=\bra{\psi}\hat{H}_{c}\ket{\psi}$. We define the approximation ratio of the QAOA as
@@
where $F_{0}$ represents the optimal value and $F_{\max}$ signifies the worst-case value.
Our study uses a QAOA with qubit numbers ranging from 3 to 12 and a depth of 1. To establish a baseline for experimental results, we present the simulation results conducted in a noise-free environment with Qiskit’s Qasm simulator in Table 1. The version of Qiskit used throughout this paper is 0.45.3. All data use 30,000 circuit repetitions (shots). We observe that the approximation ratio and success probability decrease as the number of qubits increases.
Table 1: Noise-free simulation results of the quantum approximate optimization algorithm (QAOA) for portfolio optimization using Qiskit’s Qasm simulator with 30,000 shots.
@@ II.1.2 Metric Definition
where $r_{\epsilon}$ and $p_{\epsilon}$ represent the approximation ratio and success probability obtained under noise conditions. $r_{0}$ and $p_{0}$ denote corresponding values obtained from the simulated noise-free case (Table 1). For the same problem instances, we use identical values of $r_{0}$ and $p_{0}$ for evaluation. Due to noise, the NAR and NSP typically exhibit values less than one. However, in rare instances, these metrics may exceed unity, indicating that specific noise patterns or randomness enhance performance compared to the simulated noise-free case.
@@
Without any error mitigation being applied, the NAR and NSP for algorithms executing on real noisy quantum devices are given by
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}$ $=$ $r_{b}/r_{0},$ (6)
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}$ $=$ $p_{b}/p_{0},$ (7)
@@
where B represents results obtained from real hardware without error mitigation. $r_{b}$ and $p_{b}$ represent the corresponding approximation ratio and success probability, respectively. When a specific error mitigation strategy, such as the DD sequence, is applied, the NAR and NSP are given by
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}$ $=$ $r_{d}/r_{0},$ (8)
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}$ $=$ $p_{d}/p_{0},$ (9)
@@
where DD denotes the application of DD sequences. $r_{d}$ and $p_{d}$ represent the approximation ratio and success probability, respectively, with error mitigation. To assess DD effectiveness in error mitigation, we introduce two additional metrics: $\Delta_{\mathrm{\scriptscriptstyle NAR}}$ and $\Delta_{\mathrm{\scriptscriptstyle NSP}}$. These metrics are defined as the difference between the corresponding values obtained with and without DD sequences:
$\Delta_{\mathrm{\scriptscriptstyle NAR}}$ $=$ $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}-\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}},$ (10)
$\Delta_{\mathrm{\scriptscriptstyle NSP}}$ $=$ $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}-\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}.$ (11)
@@
A positive value for $\Delta_{\mathrm{\scriptscriptstyle NAR}}$ and $\Delta_{\mathrm{\scriptscriptstyle NSP}}$ indicates a successful improvement in performance resulting from the utilization of DD sequences.
@@
We further introduce the concept of error mitigation success rate (EMSR) to quantify the robustness of an error mitigation strategy. EMSR is defined as the percentage of experimental trials where error mitigation improved the outcome compared to no mitigation. A high EMSR indicates consistent performance improvement while a low EMSR suggests limited effectiveness or potential performance degradation. We employ two EMSR metrics in this study, $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle AR}}$ and $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle SP}}$, depending on whether the approximation ratio or success probability is used. Positive values of $\Delta_{\mathrm{\scriptscriptstyle NAR}}$ and $\Delta_{\mathrm{\scriptscriptstyle NSP}}$ contribute to increased $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle AR}}$ and $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle SP}}$, respectively.
@@ II.2 Hardware Considerations
This section provides information about the IBM quantum devices used in our experiments. The QPUs with 27 qubits, namely ibmq_mumbai, ibmq_kolkata, ibm_cairo, and ibmq_ehningen, operate using basis gates {CX, ID, RZ, SX, X}, where ID represents identity gate, RZ performs a single qubit rotation around the $z$-axis, X is the NOT gate, and SX is the square root of X. On the other hand, the QPUs with 127 qubits, specifically ibm_kyoto, ibm_cusco, ibm_brisbane, and ibm_sherbrooke, utilize basis gates {ECR, ID, RZ, SX, X}, where ECR is the echoed cross-resonance gate.
Figure 1(a) illustrates the schedule of a native CX gate on ibm_cairo. This implementation employs a single-pulse gate duration of 112 dt and a cross-resonance (CR) gate duration of 544 dt, leading to a total duration of 1312 dt, where dt represents the system cycle time. In contrast, Fig. 1(b) depicts the schedule of a native ECR gate on ibm_cusco. Here, the single-pulse and CR gate durations are 88 dt and 416 dt, respectively, resulting in a shorter total duration of 920 dt. The CX gate is typically implemented using one ECR gate and multiple single-qubit gates. A further decomposition of the CX gate into ECR and single-pulse gates at pulse level enables the elimination of single-pulse gates during circuit optimization [44]. Additionally, CX-based IBM QPUs, where CX is the native gate, support the operation of CX in two directions. On the other hand, ECR-based QPUs, where ECR is the native gate, typically only support the ECR gate in one direction. Consequently, quantum circuits designed for these latter devices need to be decomposed into sequences of gates that include only the supported direction of the ECR.
@@
Figure 1: Schedule of hardware-native two-qubit gates: (a) CX gate on qubit pair (0,1) of ibm_cairo, and (b) echoed cross-resonance (ECR) gate on qubit pair (0,1) of ibm_cusco. The system cycle time ($1\mathrm{dt}$) is 2/9 ns $\approx$ 0.22 ns in ibm_cairo, while it is 0.50 ns in ibm_cusco. $D_{i}$ represents the drive channel acting on qubit $i$, and $U_{j}$ is the control channel for a corresponding qubit pair $(c,t)$ driving the control qubit $c$ at the frequency of the target qubit $t$.